OpenStudy (anonymous):

how many cubic centimeters (cm3) of olive oil have the same mass as 2.00 L of gasoline. given gasoline has a density of .66 g/mL and olive oil .92 g/mL?

OpenStudy (stamp):

What all information is given to you? What all information are you required to find? How do you go about solving for said information?

OpenStudy (stamp):

Given information:\[m_{olive\ oil}=m_{gasoline}\]\[d_{gasoline}=0.66\ g/mL\]\[d_{olive\ oil}=0.92\ g/mL\]\[V_{gasoline}=2.00\ L\]

OpenStudy (stamp):

Required information:\[Find\ V_{olive\ oil}=x\ cm^3.\]

OpenStudy (stamp):

What is your solution?

OpenStudy (anonymous):

what is given to me is olive oil has a density of .92g/mL and gasoline has a density of .66 g/mL

OpenStudy (anonymous):

i have no clue in how to solve this problem. its for chemistry. and its asking how many cm3 of olive oil have the same mass as 2.00 L of gasoline

OpenStudy (stamp):

Observe:\[m_{olive\ oil}=m_{gasoline}\]\[d=\frac{m}{V}\]\[m=d*V\]

OpenStudy (anonymous):

do i equal them to each other?

OpenStudy (stamp):

\[m_{olive\ oil}=d_{olive\ oil}*V_{olive\ oil}\]\[=(0.92\ g/mL)x\ cm^3=0.92x\] \[m_{gasoline}=d_{gasoline}*V_{gasoline}\]\[Find\ m_{gasoline}.\]Set the masses equal and solve for x.

OpenStudy (anonymous):

dont i have to convert L to grams?

OpenStudy (stamp):

You can convert L to mL or mL to L, but yes, be wary of your units. Let me know what values you obtain for m.